OpenStudy (anonymous):

Help The distance from the north pole to the equator is approximately 10,000km.Use this information to find and approximate value for the earth's radius Pic below

OpenStudy (anonymous):



OpenStudy (anonymous):

Is this a question about a sphere or about a circle?

OpenStudy (zale101):

I think its about finding Radius and circumference of circles.

OpenStudy (anonymous):

circle

OpenStudy (anonymous):

no I just have to find the radius

OpenStudy (anonymous):

Is C the centre of the circle?

OpenStudy (anonymous):

Then you use Pythagoras to find the length of the two sides (which are the same) knowing the hypotenuse. So a^2 + b^2 = c^2 a = b c = 10,000

OpenStudy (anonymous):

Or is the length of the ARC 10000? This is not a well worded question.

OpenStudy (anonymous):

If the arc is 10000, then you use the formula for circumference of a circle (2pi*r) to find the radius.

OpenStudy (anonymous):

But of course if the book is talking about the Earth, then none of this applies because the Earth is an oblate spheroid.

OpenStudy (anonymous):

I think it its its from the north pole N to the equator E

OpenStudy (anonymous):

Well use the circumference method and the pythagoras method and see which one gives you the right answer.

jhonyy9 (jhonyy9):

so the circumference of a circle is 2pir what you know that is 360 degree so what you know from the text of your exercise is that to an segment (part) with 90 degree has a circumference distance of 10000 km but you know again that this 90 degree is 1/4 of 360 so now you can calculi using this to 360/4 ----- 10000 km than to 360 ----- x ----------------------------- x = 360*10000 /(360/4) = 360*10000 * 4/360 =4*10000 =40000 km than now you know that the circumference of the Earth is 40000 km what is equale 2pir = 2*3,14*r so using this you will can easy calculi the radius 2pir=2*3,14*r = 6,28*r 6,28*r=40000 r= 40000/6,28 = 6369,4 so what mean that the radius of the earth is aproximatly 6370 km